# Waveform Optimization and Beam Focusing for Near-field Wireless Power Transfer with Dynamic Metasurface Antennas and Non-linear Energy Harvesters

### v2

@@ Abstract
Radio frequency (RF) wireless power transfer (WPT) is a promising technology for future wireless systems. However, the low power transfer efficiency (PTE) is a critical challenge for practical implementations. One of the main inefficiency sources is the power consumption and loss introduced by key components such as high-power amplifier (HPA) and rectenna, thus they must be carefully considered for PTE optimization. Herein, we consider a near-field RF-WPT system with the emerging dynamic metasurface antenna (DMA) at the transmitter and non-linear energy harvesters. We provide a mathematical framework to calculate the power consumption and harvested power from multi-tone signal transmissions. Then, we propose an approach relying on alternating optimization and successive convex approximation for waveform optimization and beam focusing to minimize power consumption while meeting energy harvesting requirements. Numerical results show that increasing the transmit tones reduces the power consumption by leveraging the rectifier’s non-linearity. Moreover, it is demonstrated that increasing the antenna length improves the performance, while both DMA and fully-digital architectures may be favorable depending on the setup. Finally, our results verify that the transmitter generates accurate energy beams pointed to devices located in the near-field, while energy beams are formed in devices’ direction in the far-field region.
@@ I Introduction
EH receivers may harvest energy from two types of sources: those that exist in the surrounding environment, and those that are specifically designated for energy transmission. From a transmission perspective, the latter is supported by wireless power transfer (WPT) technologies, e.g., based on inductive coupling, magnetic resonance coupling, laser power beaming, and radio frequency (RF) radiation. Among them, RF-WPT is promising for charging multiple users relatively far from the transmitter by exploiting the broadcast nature of wireless channels. Furthermore, this can be accomplished over the same infrastructure used for wireless communications. Notably, the most important challenge toward maturing RF-WPT is related to increasing the inherently low end-to-end power transfer efficiency (PTE) [4]. Herein, we focus on RF-WPT, which is referred to as WPT in the following.
@@ I-A Preliminaries
The end-to-end PTE depends on the performance of the key building blocks, namely energy transmitter (ET), wireless channel, and energy receiver (ER) as illustrated in Fig. 1. At first, a signal is generated and amplified using a direct current (DC) power source at the ET. Then, it is upconverted to the RF domain and transmitted over the wireless channel. Finally, the ER converts the received RF signal to DC for EH purposes. Indeed, the end-to-end PTE comprises: DC-to-RF, RF-to-RF, and RF-to-DC power conversion efficiency, i.e.,
@@
In WPT, both transmitter and receiver introduce non-linearities to the signals affecting the amount of harvested power. Indeed, an appropriately designed transmit signal leveraging these non-linearities may reduce the power consumption at the transmitter and/or increases the RF-to-DC conversion efficiency at the receiver [5, 6]. Specifically, using multiple transmit frequency tones can lead to high peak-to-average power ratio (PAPR) signals at the receiver, enhancing the rectifier RF-to-DC conversion efficiency [7]. Meanwhile, the RF signal can be focused towards the ER using energy beamforming (EB), which affects the transmit/receive waveform, to cope with the channel inefficiencies captured by $e_{2}$, thereby enhancing the amount of RF power that can be harvested [4]. Notice that the active transmit components consume power to operate, while the passive elements introduce power losses, both of which impact $e_{1}$ and must be considered. Therefore, $e_{1}$, $e_{2}$, and $e_{3}$ are correlated, suggesting that their joint optimization may lead to significant gains in terms of end-to-end PTE.
@@
The end-to-end PTE is also affected by the system power consumption. One of the main factors contributing greatly to the power consumption is the transmitter’s architecture, which also determines the beamforming approach. For example, in a fully-digital structure, each antenna element necessitates a dedicated RF chain, with its corresponding high-power amplifier (HPA), consuming a significant amount of power. An HPA aims to amplify the input signal to compensate for the path loss and fading in a wireless system. Moreover, the signal amplification in the HPA requires a DC power source, which accounts for the majority of the power consumption. The significant drawback of the fully-digital architecture is the high complexity and cost, making it impractical for applications requiring massive multiple-input multiple-output (mMIMO) implementations. Alternatively, analog architectures using, e.g., passive phase shifters, are less expensive but offer limited degrees of freedom for EB. Thus, a hybrid architecture implementation combining both approaches is often more appealing in practice. Hybrid architectures offer a trade-off between complexity (cost) and beamforming flexibility [8, 9].
@@
Although hybrid beamforming using phase shifters promotes cost reduction, it still requires complex analog networks for phase shifting. There are emerging technologies to provide hybrid beamforming capability with an even lower cost and complexity, e.g., reflective intelligent surface (RIS)-aided systems [10] and dynamic metasurface antennas (DMAs) [11]. Notice that RIS is an assisting node that provides the passive beamforming capability using reflective elements, while DMA is a transceiver consisting of configurable metamaterial elements and a limited number of RF chains. DMA avoids analog network implementation challenges and provides hybrid beamforming capability with low cost and complexity. Each of these architectures may be favorable based on the system setup. For instance, employing multiple low-cost RISs helps to cover the blind spots that are prone to weak signal reception in a large area. However, reflecting surfaces in RIS-assisted systems lack baseband processing capability to perform channel estimation and send pilot signals. Thus, acquiring accurate enough channel state information to attain a suitable passive beamforming gain might force huge overheads to the system [12]. On the other hand, DMA is a transceiver and has sufficient baseband processing capability for channel estimation. However, its implementation requires some RF chains making it more costly than RIS for large-scale implementation. All in all, both of these architectures support low-cost transmitter deployments, while the choice highly depends on the considered system setup. Interestingly, the authors in [13] utilize a system model comprising both RIS and DMA structures for uplink MIMO communication, while assuming that the channel is perfectly known.
@@ I-B Prior works
There are many works either focusing on EB, waveform optimization, or joint waveform and beamforming design for fully-digital WPT systems. The authors in [14] utilize EB to power multiple devices in a MIMO system consisting of radio stripes, while the deployment problem of this transmit architecture is investigated in [15]. Furthermore, a low-complexity beamforming design relying on the statistics of the channel is proposed in [16] to fairly power a set of single-antenna devices. In the mentioned works, none of the practical system non-linearities are considered and the focus is on the received RF power at the devices. In [17], transmit beamforming and RF and DC combining at the receiver are leveraged to increase the received DC power in a MIMO WPT system. Although this work considers the rectifier’s non-linearity, the waveform design is not investigated. Moreover, a low-complexity waveform design for single-user setups is proposed in [18], while the large-scale multi-antenna WPT scenario is addressed in [19]. The authors in [20] leverage beamforming and a multi-sine waveform in a MIMO WPT system to enhance the harvested power. Notably, the frameworks in [18, 19, 20] consider the rectifier non-linearity. Interestingly, the authors in [21] perform waveform and beamforming optimization while considering both main non-linearity sources (HPA and rectifier) aiming to maximize the harvested DC power in a WPT system.
Although most of the works on WPT systems in the literature focus on traditional fully-digital architecture, novel low-cost transmitters have also attained significant attention recently. For instance, DMA is utilized in [22, 23] for a near-field WPT system, while the authors in [24] propose a minimum-power beamforming design for meeting quality of service requirements of the users in a simultaneous wireless information and power transfer (SWIPT) system. However, none of these studies have considered the rectifier non-linearity and its impact on the harvested DC power. Notably, the joint waveform and beamforming design problem in RIS-aided WPT and SWIPT systems is investigated in [25] and [26], respectively. Furthermore, the two latter works consider the EH non-linearity at the receiver side, thus, taking into account its impact on the harvested DC power.
@@ I-C Contributions
All in all, WPT systems have received considerable attention for some time. Still, more effort is needed to reduce the system power consumption, thus increasing the end-to-end PTE. For this, low-cost multi-antenna transmitters like DMA are appealing and may pave the way for charging devices efficiently in massive IoT deployments. Moreover, multiple studies aimed to enhance the amount of harvested DC power (with rectifier non-linearity) in far-field WPT systems or received RF power (without rectifier non-linearity) in near-field WPT. As mentioned before, the receiver does not perform RF-to-DC conversion linearly and the shape of the waveform may leverage the receiver’s RF-to-DC conversion efficiency. Thus, it is imperative to take into account the impact of the rectifier’s non-linearity in the system. To the best of our knowledge, no work has yet investigated the radiative near-field power transmission and the power consumption of a multi-antenna WPT system for meeting the EH requirements of a multi-user setup while considering the receiver non-linearity, especially when using low-cost transmitters. Herein, we aim precisely to fill this research gap. Our main contributions are as follows:
@@
1) We formulate a joint waveform optimization and beam focusing problem for a multi-user multi-antenna WPT system with both a fully-digital and a DMA architecture. Due to the huge potential of near-field WPT systems for future practical WPT applications [3], we present our system model relying on a near-field wireless channel, which can inherently capture far-field conditions as well. Notice that there are some previous works [25, 21], which focused on increasing the amount of harvested power in far-field WPT systems while considering the receiver non-linearity. However, the literature lacks a minimum-power waveform and beamforming design (even for far-field channels), which can deal with meeting the EH requirements. This is a critical gap to fill since such formulation mimics a practical setup where the EH users inform their DC power demands and the WPT system must serve them with minimum power consumption, thus maximum end-to-end PTE. Since the HPA is an active element incurring most of the power consumption at the transmitter side, we model the power consumption of a class-B HPA as a function of its output power. Notably, our problem for fully-digital architecture shares some similarities with the one discussed in [21] as objectives and constraints are interchanged. However, the main focus of our work here is on the DMA-assisted system, which introduces much more complexity to the problem due to the coupling between the variables and their Lorentzian-constrained phase response. Note that the phase shift introduced by the metamaterial elements is correlated with their amplitude, which results in a different beamforming problem than other architectures, e.g., RIS-assisted systems [26] and phase shifter-based hybrid beamforming [9]. Mathematically, when dealing with those latter architectures, both RIS passive elements or phase shifters introduce a phase shift to the signal with constant loss [^1] Note that in most of the works in the literature, without loss of generality, this phase shifting process by the analog network or RIS elements is considered to be lossless., while each phase-shifting configuration in DMA elements leads to a different propagation loss introduced to the transmit signal. The complexities associated with our specific problem make the existing optimization frameworks for WPT systems inapplicable to our system calling for novel approaches.
@@
2) We propose a method relying on alternating optimization and successive convex approximation (SCA) to efficiently solve the waveform and beamforming optimization problem in the DMA-assisted WPT system. Specifically, we decouple the optimization problem to maximize the minimum received DC power by tuning the frequency response of the metamaterial elements, while minimizing the consumed power for meeting the EH requirements when optimizing the digital precoders. Generally, a huge complexity is introduced to the waveform optimization problems by time sampling since the number of samples should be relatively large to result in a reliable framework [21, 6, 20]. To cope with this, we reformulate the received DC power of the users based on the spectrum of the received waveform, which removes the time dependency in the problem. Then, the metamaterial elements and the digital precoders are alternatively optimized using SCA. Motivated by the influence of variable initialization on the SCA performance, we propose a low-complexity initialization algorithm for the digital precoders and DMA weights by leveraging the channel characteristics and dedicating RF signals to the ERs. Furthermore, the complexity of the proposed optimization framework scales with the number of users, antenna elements, and frequency tones.
@@
3) We illustrate the convergence of the proposed optimization method numerically and show that the complexity increases with the antenna length, number of transmit tones, and number of devices. Furthermore, we provide evidence that increasing the antenna length and the number of transmit tones reduces power consumption, while it increases with the number of devices and user distance. Moreover, our findings evince that DMA performs better in terms of power consumption when the number of RF chains and transmit tones are relatively low, while the fully-digital architecture becomes favorable when the mentioned parameters are sufficiently large. This also depends on the specific HPAs’ saturation power, number of devices, and user distance. Additionally, we verify by simulation that the transmitter can accurately focus the energy on the receiver location in the near-field region, while energy beams are only formed toward specific directions in the far-field region.
@@ II System Model
We consider a multi-antenna WPT system to charge $M$ single-antenna EH devices. The received RF power at the ER is transferred into the rectifier input using a matching network. Then, it is converted to DC power by the rectifier, while $\bar{P}_{m}$ denotes the EH requirement of the $m$th ER.

### v3

@@ Abstract
Radio frequency (RF) wireless power transfer (WPT) is a promising technology for future wireless systems. However, the low power transfer efficiency (PTE) is a critical challenge for practical implementations. One of the main inefficiency sources is the power consumption and loss introduced by key components such as high-power amplifier (HPA) and rectenna, thus they must be carefully considered for PTE optimization. Herein, we consider a near-field RF-WPT system with a dynamic metasurface antenna (DMA) at the transmitter and non-linear energy harvesters. We provide a mathematical framework to calculate the power consumption and harvested power from multi-tone signal transmissions. Based on this, we propose an approach relying on alternating optimization and successive convex approximation for waveform optimization and beam focusing to minimize power consumption while meeting energy harvesting requirements. Numerical results show that increasing the number of transmit tones reduces the power consumption by leveraging the rectifier’s non-linearity more efficiently. Moreover, they demonstrate that increasing the antenna length improves the performance, while DMA outperforms fully-digital architecture in terms of power consumption. Finally, our results verify that the transmitter focuses the energy on receivers located in the near-field, while energy beams are formed in the receivers’ direction in the far-field region.
@@ I Introduction
EH receivers may harvest energy from two types of sources: those that exist in the surrounding environment, and those that are specifically designated for energy transmission. From a transmission perspective, the latter is supported by wireless power transfer (WPT) technologies, e.g., based on inductive coupling, magnetic resonance coupling, laser power beaming, and radio frequency (RF) radiation. Among them, RF-WPT is promising for charging multiple receivers relatively far from the transmitter by exploiting the broadcast nature of wireless channels. Furthermore, this can be accomplished over the same infrastructure used for wireless communications. Notably, the most important challenge toward maturing RF-WPT is related to increasing the inherently low end-to-end power transfer efficiency (PTE) [4]. Herein, we focus on RF-WPT, which is referred to as WPT in the following.
@@ I-A Preliminaries
The end-to-end PTE depends on the performance of the key building blocks of a WPT system, namely energy transmitter (ET), wireless channel, and energy receiver (ER) as illustrated in Fig. 1. At first, a signal is generated and amplified using a direct current (DC) power source at the ET. Then, it is upconverted to the RF domain and transmitted over the wireless channel. Finally, the ER converts the received RF signal to DC for EH purposes. Indeed, the end-to-end PTE comprises: DC-to-RF, RF-to-RF, and RF-to-DC power conversion efficiency, i.e.,
@@
In WPT, both transmitter and ER introduce non-linearities to the signals affecting the amount of harvested power. Indeed, an appropriately designed transmit signal leveraging these non-linearities may reduce the power consumption at the transmitter and/or increase the RF-to-DC conversion efficiency at the ER [5, 6]. Specifically, using multiple transmit frequency tones can lead to high peak-to-average power ratio (PAPR) signals at the ER, enhancing the rectifier RF-to-DC conversion efficiency [7]. Meanwhile, the RF signal can be focused towards the ER using energy beamforming (EB), which affects the transmit/receive waveform, to cope with the channel inefficiencies captured by $e_{2}$, thereby enhancing the amount of RF power that can be harvested [4]. Notice that the active transmit components consume power, while the passive elements introduce power losses, both of which impact $e_{1}$ and must be considered. Therefore, $e_{1}$, $e_{2}$, and $e_{3}$ are correlated, suggesting that their joint optimization may lead to significant gains in terms of end-to-end PTE.
@@
The end-to-end PTE is also affected by the system power consumption. One of the main factors contributing greatly to the power consumption is the transmitter’s architecture, which also determines the beamforming approach. For example, in a fully-digital structure, each antenna element necessitates a dedicated RF chain with its corresponding high-power amplifier (HPA), consuming a significant amount of power. An HPA aims to amplify the input signal to compensate for the path loss and fading in a wireless system. Moreover, the signal amplification in the HPA requires a DC power source, which accounts for the majority of the power consumption. The main drawback of the fully-digital architecture is the high complexity and cost, making it impractical for applications requiring massive multiple-input multiple-output (MIMO) implementations. Alternatively, analog architectures using, e.g., passive phase shifters, are cheaper but offer limited degrees of freedom for EB. Thus, a hybrid architecture implementation combining both approaches is often more appealing in practice. Hybrid architectures offer a trade-off between complexity (cost) and beamforming flexibility [8, 9].
@@
Although hybrid beamforming using phase shifters promotes cost reduction, it still requires complex analog networks for phase shifting. Moreover, the complexity of traditional phase-shifter-based hybrid architectures scales with the complexity of this analog network. For instance, a fully-connected network of phase shifters for connecting the output of the RF chains to the antenna elements is the most complex architecture, while there are some other low-complexity architectures, e.g., dynamic array-of-subarrays architecture [10, 11]. There are emerging technologies to provide hybrid beamforming capability with an even lower cost and complexity, e.g., reconfigurable intelligent surface (RIS)-aided systems [12] and dynamic metasurface antennas (DMAs) [13]. Notice that, RIS is an assisting node that provides the passive beamforming capability using reflective elements, [^1] There are also some novel RIS-based transmit structures that leverage RIS with active elements to provide extra capabilities [14]. while DMA is a transceiver consisting of configurable metamaterial elements and a limited number of RF chains. DMA avoids analog network implementation challenges and provides hybrid beamforming capability with low cost and complexity. Each of these architectures may be preferred based on the system setup. For instance, employing multiple low-cost RISs helps to cover the blind spots that are prone to weak signal reception in a large area. However, reflecting surfaces in RIS-assisted systems lack baseband processing capability to perform channel estimation and send pilot signals. Thus, acquiring accurate enough channel state information (CSI) to attain a suitable passive beamforming gain might force huge overheads to the system [15]. On the other hand, DMA is a transceiver and has sufficient baseband processing capability for channel estimation. However, its implementation requires some RF chains making it more costly than RIS for large-scale implementation. All in all, both of these architectures support low-cost transmitter deployments, while the choice highly depends on the considered system setup. Interestingly, the authors in [16] utilize a system model comprising both RIS and DMA structures for uplink MIMO communication while assuming that the channel is perfectly known.
@@ I-B Prior works
There are many works either focusing on EB, waveform optimization, or joint waveform and beamforming design for fully-digital WPT systems. The authors in [17] utilize EB to power multiple devices in a MIMO system consisting of radio stripes, while the deployment problem of this transmit architecture is investigated in [18]. Furthermore, a low-complexity beamforming design relying on the statistics of the channel is proposed in [19] to fairly power a set of single-antenna devices. In the mentioned works, none of the practical system non-linearities are considered and the focus is on the received RF power at the devices. In [20], transmit beamforming and RF and DC combining at the ER are leveraged to increase the received DC power in a MIMO WPT system. Although this work considers the rectifier’s non-linearity, the waveform design is not investigated. Moreover, a low-complexity waveform design for single-user setups is proposed in [21], while the large-scale multi-antenna WPT scenario is addressed in [22]. The authors in [23] leverage beamforming and a multi-sine waveform in a MIMO WPT system to enhance the harvested power. Notably, the frameworks in [21, 22, 23] consider the rectifier non-linearity. Interestingly, the authors in [24] perform waveform and beamforming optimization while considering both main non-linearity sources (i.e., HPA and rectifier) aiming to maximize the harvested DC power in a WPT system.
Although most of the works on WPT systems in the literature focus on a traditional fully-digital architecture, novel low-cost transmitters have also attained significant attention recently. For instance, DMA is utilized in [25, 26] for a near-field WPT system, while the authors in [27] propose a minimum-power beamforming design for meeting quality of service requirements of the ERs in a simultaneous wireless information and power transfer (SWIPT) system. However, none of these studies have considered the rectifier non-linearity and its impact on the harvested DC power. Notably, the joint waveform and beamforming design problem in RIS-aided WPT and SWIPT systems is investigated in [28] and [29], respectively. Furthermore, the two latter works consider the EH non-linearity at the ER side, thus, taking into account its impact on the harvested DC power.
@@ I-C Contributions
All in all, WPT systems have received considerable attention for some time. Still, more effort is needed to reduce the system power consumption, thus increasing the end-to-end PTE. For this, low-cost multi-antenna transmitters like DMA are appealing and may pave the way for charging devices efficiently in massive IoT deployments. Moreover, multiple studies aimed to enhance the amount of harvested DC power (with rectifier non-linearity) in far-field WPT systems or received RF power (without rectifier non-linearity) in near-field WPT. As mentioned before, the ER does not perform RF-to-DC conversion linearly and a proper waveform design must leverage the ER’s RF-to-DC conversion efficiency dynamics. Thus, it is imperative to take into account the impact of the rectifier’s non-linearity. To the best of our knowledge, no work has yet investigated the power consumption of a multi-antenna WPT system for meeting the EH requirements of a multi-user setup while considering the ER non-linearity, especially when using low-cost transmitters. Herein, we aim precisely to fill this research gap. Our main contributions are as follows:
@@
1) We formulate a joint waveform optimization and beam focusing problem for a multi-user multi-antenna WPT system with both a fully-digital and a DMA architecture. Due to the huge potential of near-field WPT systems for future practical WPT applications [3], we present our system model relying on a near-field wireless channel, which can inherently capture far-field conditions as well. Notice that there are some previous works [28, 24] focused on increasing the amount of harvested power in far-field WPT systems while considering the ER non-linearity. However, the literature lacks a minimum-power waveform and beamforming design (even for far-field channels) that can deal with meeting EH requirements. This is a critical gap to fill since such formulation mimics a practical setup where the EH receivers inform their DC power demands and the WPT system must serve them with minimum power consumption, thus maximum end-to-end PTE. Since the HPA is an active element incurring most of the power consumption at the transmitter side, we model the power consumption of a class-B HPA as a function of its output power. Notably, our problem for fully-digital architecture shares some similarities with the one discussed in [24] as objectives and constraints are interchanged. However, there are two main differences. First, since the problem is highly non-linear and non-convex, aiming for a minimum-power waveform and beamforming design leads to a different and complex problem due to the appearance of the highly non-convex constraints and a non-convex power consumption term relying on the transmit waveform. Second, the main focus of our work here is on a DMA-assisted system, which introduces much more complexity to the problem due to the coupling between the optimization variables, i.e., the frequency response of the elements and the digital beamforming complex weights, and the Lorentzian-constrained phase response of the elements. Note that the phase shift introduced by the metamaterial elements is correlated with their amplitude, which results in a different beamforming problem than other architectures, e.g., RIS-assisted systems [29] and phase shifter-based hybrid beamforming [9]. Mathematically, when dealing with those latter architectures, both RIS passive elements or phase shifters introduce a phase shift to the signal with constant loss, [^2] Note that in most of the works in the literature [29, 27], without loss of generality, this phase shifting process by the analog network or RIS elements is considered to be lossless. while each phase-shifting configuration in DMA elements leads to a different propagation loss introduced to the transmit signal. Meanwhile, observe that our utilized DMA model is similar to [25, 30], which is the standard model in the literature [13]. However, we investigate this model for WPT when considering ER non-linearity, which leads to a different problem in terms of complexity and required solutions, especially when aiming for minimum-power waveform and beamforming design instead of aiming for received power maximization. All in all, the complexities associated with our specific problem make the existing optimization frameworks for WPT systems inapplicable to our system, calling for novel approaches.
@@
2) We propose a method relying on alternating optimization and successive convex approximation (SCA) to efficiently solve the waveform and beamforming optimization problem in the DMA-assisted WPT system. Specifically, we decouple the optimization problem to maximize the minimum received DC power by tuning the frequency response of the metamaterial elements, while minimizing the power consumption needed for meeting the EH requirements when optimizing the digital precoders. Generally, a huge complexity is introduced to the waveform optimization problems by time sampling since the number of samples should be relatively large to result in a reliable framework [24, 6, 23]. To cope with this, we reformulate the received DC power of the ERs based on the spectrum of the received waveform, which removes the time dependency in the problem. Then, the metamaterial elements and the digital precoders are alternatively optimized using SCA. Motivated by the influence of variable initialization on the SCA performance, we propose a low-complexity initialization algorithm for the digital precoders and DMA weights by leveraging the channel characteristics and dedicating RF signals to the ERs. Furthermore, the complexity of the proposed optimization framework scales with the number of ERs, antenna elements, and frequency tones. Note that the location of the ERs is considered to be perfectly known at the transmitter side, leading to perfect CSI at the transmitter. Indeed, the proposed framework can be used for any channel model as long as the CSI is known. However, the focus of the work is on the radiative near-field region for WPT applications; thus, we also use the term beam focusing in addition to traditional beamforming, which is typical for far-field communications.
@@
3) We illustrate the convergence of the proposed optimization method numerically and show that the complexity increases with the antenna length, number of transmit tones, and number of receivers. Furthermore, we provide evidence that increasing the antenna length and the number of transmit tones reduces power consumption, while it increases with the number of receivers and ER distance. Moreover, our findings evince that the DMA-assisted architecture outperforms the fully-digital counterpart in terms of power consumption in our system. Moreover, the performance gap between these architectures is shown to depend on the specific HPAs’ saturation power, number of devices, and ER distance. Additionally, we verify by simulation that the transmitter can accurately focus the energy on the ER location in the near-field region, while energy beams are only formed toward specific directions in the far-field region.
@@ II System Model
Figure 2: Transmit antenna architectures. (a) fully-digital architecture (left) and (b) DMA-assisted architecture (right).
We consider a multi-antenna WPT system to charge $M$ single-antenna EH devices. The received RF power at the ER is transferred to the rectifier input using a matching network. Then, it is converted to DC power by the rectifier, while $\bar{P}_{m}$ denotes the EH requirement of the $m$th ER.